OpenStudy (anonymous):

A long distance trucker traveled 260 miles in one direction during a storm. The return trip in rainy weather was accomplished at double the speed and took 5 hours less time. Find the speed going.

OpenStudy (anonymous):

let x be speed going then x = 260 / t..................(1) speed returning = 2x and time = t-5 hours 2x = 260 / t-5....................(2) multiply (1) by 2: 2x = 520 /t so 520/t = 260 / t-5 520t - 2600 = 260t 260t = 2600 t = 10 speed x = 260 / 10 = 26 mph

OpenStudy (anonymous):

THX!!

OpenStudy (anonymous):

no probs